In Bohr’s model of the hydrogen atom, the electron travels with speed 2.2 x 106m/s in a circle (r= 5.3 x 10-11m) about the nucleus. Find the value of B at the nucleus due to the electrons motion. Assume vacuum.


1
Expert's answer
2021-04-26T18:31:47-0400

To be given in question

Speed of electron (V)="2.2\times10^6 meter/sec"

Radius of circle (r)="5.3\times10^{-11}meter"

Mass of electron"(m)" ="9.1\times10^{-31} kg"

Charge of electron"(q)" ="1.6\times20^{-19}C"

To be asked in question

Magnetic field (B)=?

We know that

Magnetic force=centripetal force

"qvB =\frac{mv^2}{r} \rightarrow(1)"

"B=\frac{mv}{qr}\rightarrow(2)"

Put value in eqution (2)

"B=\frac{9.1\times10^{-31}\times2.2\times10^6}{1.6\times10^{-19}\times5.3\times10^{-11}}"

"B=236084.9057 tesla"

"B=2.36084\times10^5 T"